# version 0.6

### v3

@@ 6.2.4 Disturbance Action Controllers
We next define the most important set of policies in nonstochastic control, known as the disturbance action controllers, or DAC. The main advantages of DAC as opposed to GLC, LDC and linear controllers are that the cost of control is convex in their parametrization, and thus allows for provably efficient optimization. Moreover, this set of policies is stronger than GLC in the sense that it $\varepsilon$-approximates them, as we shortly show.
@@
A disturbance action policy $\pi_{K_{1:T},M_{1:h}}$ is parameterized by the matrices $M_{0:h}=[M_{0},{M}_{1},\ldots,{M}_{h}]$ and stabilizing controllers $\{K_{t}\}$. It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
@@
Denote by $\Pi^{DAC}_{h,\gamma}$ the set of all disturbance action policies with $K_{1:T}$ being $\gamma$-stabilizing linear controllers for the system and
@@
Any DAC policy in $\Pi_{h,\gamma}^{DAC}$ is $\gamma^{2}$-stabilizing.
@@
Next, since $K_{t}$’s are stabilizing, we can bound the historically distant terms as below. Observe that the state at time $t$, following any DAC policy $\pi$, is given by
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$
$=A_{t}\mathbf{x}_{t}+B_{t}(K_{t}\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i})+\mathbf{w}_{t}$
$=(A_{t}+B_{t}K_{t})\mathbf{x}_{t}+B_{t}\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ using $\text{using }M_{0}=I$
$=\sum_{j=0}^{t-1}\left(\left[\prod_{\tau=t-j}^{t}(A_{\tau}+B_{\tau}K_{\tau})\right]B_{t-j}\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}\right)$
$=\sum_{j=0}^{t-1}\left(\left[\prod_{\tau=t-j}^{t}(A_{\tau}+B_{\tau}K_{\tau})\right]B_{t-j}\tilde{\mathbf{w}}_{t-j}\right).$
@@
One of the main appeals of the DAC class is the following approximation lemma which shows that for time-invariant systems DAC are in fact a stronger class than GLC, and thus LDC. This is a strong guarantee, since it implies that competing with the best policy in DAC implies competing with the best GLC/LDC.
We start by showing that DAC approximate all linear controllers. This is not a vacuous statement, since the $K_{t}$ terms for DAC is now only assumed to be stabilizing, and not necessarily optimal for the actual sequence.
@@
Consider a time-invariant linear dynamical system $\{A,B\}$. The class $\Pi^{DAC}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{LIN}_{\gamma^{\prime}}$ for $h=\Omega(\gamma^{\prime}\log(\frac{\gamma^{\prime}}{\varepsilon}))$ and $\gamma=\gamma^{\prime 2}$.
@@
by using the $\gamma$-stability of $(A,B)$. The approximation in terms of states and controls implies $\varepsilon$-approximation of policy class for any Lipchitz cost sequence, concluding the proof.
Recall that GLC can be lifted to a linear dynamical system. Thus a corollary of the above lemma is as follows, and is left as Exercise 6.4.
Corollary 6.11.
Consider a time invariant linear dynamical system given by $\{A,B\}$. The class $\Pi^{DAC}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{GLC}_{h^{\prime},\gamma^{\prime}}$, for parameters $h=O(h^{\prime}+\gamma^{\prime}\log\frac{\gamma^{\prime}}{\varepsilon})$, $\gamma=O(\gamma^{\prime 2})$.
@@ 6.4 Exercises
6.2. Prove Lemma 6.6.
6.3. Prove Lemma 6.10.
6.4. Prove Corollary 6.11.
@@
In this chapter we consider the main new directions that have been taken in the machine learning literature for the control problem.
These new developments are a major departure from classical control, and in recent years have produced methods that compete favorably and many times surpass classical methods. A few characteristics of the new approaches are listed here:
@@
    Instead of pre-computing an optimal policy, online control seeks to learn a policy using online learning techniques. This fits the framework known as “adaptive control”.
@@
    In contrast to adaptive control, online control optimizes a rigorous performance metric of regret vs. the best policy in hindsight. Online and non-stochastic control algorithms have provable finite-time regret guarantees.
@@
    In online nonstochastic control, the loss functions can be general Lipschitz convex losses that are chosen adversarially, as opposed to known quadratic losses in classical control.
@@
    Online and non-stochastic control deploy improper learning by convex relaxation: this means they compete with a certain policy class using a different policy class. This difference in learned policies allows us to side-step the computational hardness results we have previously explored, and obtain rigorous guarantees for efficient algorithms.
@@ 7.1 From Optimal and Robust to Online Control
Recall that the problem of optimal control of a linear time-invariant dynamical system (LDS) is given by the equations
$\min_{\mathbf{u}(\mathbf{x})}$ $\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$
subject to $\text{subject to }\mathbf{x}_{t+1}$ $={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},$
where $\mathbf{x}_{t}$ is the state of the system, $\mathbf{u}_{t}$ is the control input, and $\mathbf{w}_{t}$ is the perturbation at time $t$.
Previous chapters showed the optimal control policy to be a linear function of the state when we assumed $\left\{\mathbf{w}_{t}\right\}_{t\in[T]}$ to be i.i.d. zero mean and bounded variance perturbations, and $c_{t}$ to be quadratic loss functions. Specifically, for positive semi-definite matrices ${Q}$ and ${R}$, the mathematical control formulation becomes
$\min_{\mathbf{u}(\mathbf{x})}$ $\sum_{t=1}^{T}\mathbf{x}_{t}^{T}{Q}\mathbf{x}_{t}+\mathbf{u}_{t}^{T}{R}\mathbf{u}_{t}$
subject to $\text{subject to }\mathbf{x}_{t+1}$ $={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},\quad\mbox{ }\mathbf{w}_{t}\sim\mathcal{D},\mbox{ }\mathbb{E}\mathbf{w}_{t}=\mathbf{0}.$
We have shown in chapter 5 that the optimal solution to this formulation is given by the policy
$\mathbf{u}^{\ast}_{t}={K}\mathbf{x}_{t},$
where ${K}$ depends on ${A}$, ${B}$, ${Q}$ and ${R}$. The proof follows by inductively solving the Bellman optimality equation at each step $t$, starting from time step $T$ and moving backwards to time step $1$.
However, this optimal solution requires two conditions:
@@
    The loss function needs to be a quadratic function.
@@
    The perturbations are i.i.d. across time (with additional constraints on the mean and the variance).
These conditions are in fact necessary for solving the Bellman optimality equation. Thus, this approach is not as general as we would like. We are missing out on other loss functions commonly used in practice, e.g. cross entropy, or loss functions with regularization terms.
Furthermore and more importantly, many times the perturbations incurred in practice are not i.i.d., in which case we lose all provable guarantees!
An alternative to optimal control is robust, or $H_{\infty}$ control theory. In this formulation, the goal is to optimize the control for an adversarial noise in a minimax sense, as per the mathematical formulation:
$\min_{\mathbf{u}(\mathbf{x})}\max_{\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{1:T}\mathclose{\|}\leq 1}$ $\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$
subject to $\text{subject to }\mathbf{x}_{t+1}$ $={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
This approach also suffers from several downsides:
  1. 1.
    The above problem is computationally ill-behaved for non-quadratic loss functions.
  2. 2.
    It is non-adaptive in nature, i.e. the system is too pessimistic for cases when the perturbations are well-behaved, like i.i.d. Gaussian perturbations.
The natural question at this point is: Does there exist an approach that gives the best of both robust and optimal control, and generalizes to any convex cost function? That’s exactly non-stochastic control theory which we define next. But first, we present a motivating example.
7.1.1 Motivating Example for Online Control
Consider the problem of flying a drone from source to destination subject to unknown weather conditions. The aerodynamics of flight can be modeled sufficiently well by time-varying linear dynamical systems, and existing techniques are perfectly capable of doing a great job for indoor flight. However, the wind conditions, rain and other uncertainties are a different story. Certainly the wind is not an i.i.d. Gaussian random variable! Optimal control theory, which exactly assumes this zero-mean noise, is therefore overly optimistic and inaccurate for our problem.
We would like an adaptive control theory which allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuver when this is required by the conditions. Can we design a control theory that will allow us to take into account the specific instance perturbations and miss-specifications, and give us finite time provable guarantees? This is the subject of online non-stochastic control!
@@ 7.2 The Online Nonstochastic Control Problem
When dealing with non-stochastic, a.k.a. arbitrary or even adversarial perturbations, the optimal policy is not a-priory clear. Rather, an optimal policy for the observed perturbations can be determined only in hindsight.
This is the reason that non-stochastic control shifts the performance metric from optimality, to regret vs. the best policy in hindsight from a certain reference class. We have already defined regret in full generality in Definition 1.3, and in this section give a more detailed treatment that specializes to time varying linear dynmical systems.
At this point the reader may wonder: Why should we compare to a reference class, as opposed to all possible policies? There are two answers to this question:
First, the best policy in hindsight may be very complicated to describe and reason about. This argument was made by the mathematician Tyrell Rockafellar; see the bibliographic section for more details. Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret vs. the best policy in hindsight. Rather, a different performance metric called competitive ratio can be analyzed. This will be surveyed later in the book.
We now turn to define policy regret with respect to a reference class of policies. In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$, and then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$, according to an adversarially chosen loss function. Let $\Pi=\{\pi:\mathbf{u}\mapsto\mathbf{x}\}$ be a class of policies. The regret of the controller w.r.t. $\Pi$ is defined as follows.
@@
Let $\{A_{t},B_{t}\}$ be a time varying LDS. The regret of an online control algorithm over $T$ iterations w.r.t. the class of policies $\Pi$ is given by:
@@
$\mathbf{u}_{t}^{\pi}$ $=\pi(\mathbf{x}_{t}^{\pi})$
$\mathbf{x}_{t+1}^{\pi}$ $=A_{t}\mathbf{x}_{t}^{\pi}+B_{t}\mathbf{u}_{t}^{\pi}+\mathbf{w}_{t}.$
Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers.
The non-stochastic control problem can now be stated as the quest to find an efficient algorithm that minimizes the worst-case regret vs. meaningful policies classes that we examine henceforth. It is important to mention that the algorithm does not have to belong to the comparator set of policies $\Pi$! Indeed, in the most powerful results, we will see that the algorithm will learn a policy that is strictly not contained in the comparator class.

### v4

@@ 6.3.2 Disturbance Action Controllers
We next define the most important set of policies in online nonstochastic control, known as the disturbance action controllers, or DAC. The main advantages of DAC as opposed to linear controllers, are that the cost of control is convex in their parametrization, and thus allows for provably efficient optimization. Moreover, this set of policies is stronger than linear policies in the sense that it $\varepsilon$-approximates them, as we show below.
@@
A (linear) disturbance action policy $\pi_{K_{1:T},M_{1:h}}$ is parameterized by the matrices $M_{0:h}=[M_{0},{M}_{1},\ldots,{M}_{h}]$ and stabilizing controllers $\{K_{t}\}$. It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
@@
Denote by $\Pi^{D}_{h,\gamma}$ the set of all disturbance action policies with $K_{1:T}$ being $\gamma$-stabilizing linear controllers for the system and
@@
Any DAC policy in $\Pi_{h,\gamma}^{D}$ is $\gamma^{2}$-stabilizing.
@@
Next, since $K_{t}$’s are stabilizing, we can bound the historically distant terms as below. For simplicity assume $K_{t}=K$ are all equal. Observe that the state at time $t$, following any DAC policy $\pi$, is given by
$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}$
$=A\mathbf{x}_{t}+B(K\mathbf{x}_{t}+\sum_{i=1}^{h}M_{i}\mathbf{w}_{t-i})+\mathbf{w}_{t}$
$=(A+BK)\mathbf{x}_{t}+B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i}$ using $\text{using }M_{0}=I$
$=\sum_{j=0}^{t}\left((A+BK)^{j}B\sum_{i=0}^{h}M_{i}\mathbf{w}_{t-i-j}\right)$
$=\sum_{j=0}^{t-1}\left((A+BK)^{j}B\tilde{\mathbf{w}}_{t-j}\right).$
@@
One of the main appeals of the DAC class is the following approximation lemma which shows that for time-invariant systems DAC are in fact a stronger class than linear controllers. Notice that the $K_{t}$ terms for DAC are only assumed to be stabilizing, and not necessarily optimal for the actual sequence.
@@
Consider a time-invariant linear dynamical system $\{A,B\}$. The class $\Pi^{D}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{L}_{\gamma^{\prime}}$ for $h=\Omega(\gamma^{\prime}\log(\frac{\gamma^{\prime}}{\varepsilon}))$ and $\gamma=\gamma^{\prime 2}$.
@@
by using the $\gamma$-stability of $(A,B)$. The approximation in terms of states and controls implies $\varepsilon$-approximation of policy class for any Lipschitz cost sequence, concluding the proof.
@@ 6.4 Bibliographic Remarks
Regret minimization, online learning, and online convex optimization have a rich history and have been studied for decades in the machine learning and game theory literature. The reader is referred to the comprehensive texts of [CBL06, Haz16] for a detailed treatment.
@@
The study of regret in the context of control has roots going back to the work of [AYS11, AYBK14]. These works, as well as many that followed [CHK+18, AHS19], study regret in a stochastic setting, where the perturbations are stochastic. In such settings, optimal control is well defined. The full power of regret analysis for control, involving both adversarial perturbations, adversarial loss functions, and computational considerations, was initiated in [ABH+19].
@@ 6.5 Exercises
6.1. In this exercise we prove that the optimal solution to a linear control problem can be complex to describe.
  1. 1.
    Consider a linear dynamical system in one dimension, and loss functions of the form $f_{t}(\mathbf{x})=\max\{0,\mathbf{x}-c\}$, for $c_{t}\in{\mathbb{R}}$. Show that the optimal policy in hindsight can have as many domains as loss functions.
  2. 2.
    Consider a similar setting in higher dimensions, and give a bound on the number of domains as a function of the number of loss functions and the dimension.
6.2. Consider the setting of time varying linear dynamical systems, i.e.
$\mathbf{x}_{t+1}=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
Prove that any DAC according to definition 6.5 is stable, i.e. prove the analogue of Lemma 6.6 for time varying systems.
@@
In this chapter, we explore recent advancements in machine learning that have significantly reshaped the control problem. These developments mark a major departure from classical control theory and, in many cases, have led to methods that not only compete with but often outperform traditional approaches. The key characteristics of these modern techniques are as follows:
@@
    Online Learning Over Precomputed Policies Unlike classical control, which relies on precomputing an optimal policy, online control adopts adaptive learning techniques to update policies dynamically. This aligns with the broader framework of adaptive control, where policies evolve in response to observed system behavior.
@@
    Regret Minimization as a Performance Metric While adaptive control focuses on learning a stabilizing policy, online control introduces a rigorous performance metric: regret, defined as the difference in cost between the learned policy and the best policy in hindsight. Unlike classical approaches, online nonstochastic control algorithms provide finite-time regret guarantees, ensuring competitive performance even under adversarial conditions.
@@
    General Loss Functions and Adversarial Perturbations Classical control typically assumes quadratic cost functions with known structures. In contrast, online nonstochastic control allows for general convex loss functions, which may be selected adversarially. This greater flexibility enables controllers to adapt to more diverse and complex real-world scenarios.
@@
    Improper Learning via Convex Relaxation A fundamental breakthrough in online nonstochastic control is the use of improper learning, where the algorithm competes against a given policy class while learning from a broader class of policies. By leveraging convex relaxation techniques, these methods bypass computational hardness barriers that traditionally limit optimal control, allowing for efficient algorithms with strong theoretical guarantees.
These innovations redefine how we approach control problems, shifting the focus from static policy optimization to adaptive, regret-minimizing strategies that operate effectively in dynamic and adversarial environments.
@@ 7.1 From Optimal and Robust to Online Control
Classical control theory has been primarily shaped by two paradigms: optimal control and robust control. These approaches have been instrumental in various engineering applications, but they rely on specific assumptions about disturbances and system dynamics.
  • Optimal control assumes that disturbances follow a known stochastic model, often Gaussian. The objective is to design a control policy that minimizes the expected cumulative cost. Notable methods such as the Linear Quadratic Regulator (LQR) and stochastic dynamic programming fall into this category.
  • Robust control, in contrast, takes a worst-case perspective. It ensures stability and performance even under the most adverse conditions, assuming perturbations belong to a predefined uncertainty set. A key example is H-infinity control.
While these frameworks have been successful, real-world control tasks often involve uncertainties that do not fit neatly into either a probabilistic model or a worst-case adversarial setting. Consider the challenge of controlling a drone in an unpredictable environment:
  • A stochastic model of wind disturbances may be unrealistic since actual wind patterns can exhibit long-term correlations or abrupt shifts.
  • A worst-case robust approach may result in excessively conservative control, leading to inefficient and overly cautious maneuvers.
These limitations motivate the need for a more adaptive approach—one that can learn and adjust dynamically based on observed disturbances rather than relying on rigid assumptions. This leads us to online nonstochastic control, a fundamentally different paradigm where controllers adapt in real-time without assuming a fixed model for disturbances.
7.1.1 A Shift to Online Nonstochastic Control
Rather than relying on predefined stochastic or worst-case assumptions, online nonstochastic control introduces a new approach based on sequential decision-making and regret minimization. The core principles are:
@@ 7.1.1 A Shift to Online Nonstochastic Control
    No Assumption on the Noise Model – Disturbances (e.g., wind forces, external perturbations) are treated as an arbitrary sequence, potentially chosen adversarially. The controller must adapt dynamically rather than relying on a fixed distribution.
@@
    Learning and Adaptation Over Time – Instead of precomputing an optimal policy, online control algorithms iteratively refine their strategies using techniques from online convex optimization and sequential decision-making.
  3. 3.
    Regret Minimization as the Performance Metric – Unlike optimal control, which minimizes expected cost, or robust control, which prepares for the worst case, online control seeks to minimize regret. Regret quantifies how much worse the controller performs compared to the best policy in hindsight.
To build intuition, consider the problem of controlling an autonomous vehicle in an unknown city. If the vehicle had complete foresight of future traffic conditions, it could compute the optimal route in advance. However, without such knowledge, the vehicle must continuously adjust its route based on real-time observations. Regret measures how far the vehicle’s performance deviates from the best possible path that could have been chosen in hindsight.
This shift in perspective enables online controllers to handle a wide range of real-world disturbances, achieving adaptability, robustness, and computational efficiency without requiring strong assumptions about uncertainty.
In the following sections, we formally define the online nonstochastic control problem and introduce key algorithms that provide provable performance guarantees.
@@ 7.2 The Online Nonstochastic Control Problem
The shift from classical control to online nonstochastic control requires a fundamental rethinking of how we handle uncertainty. Unlike optimal control, which assumes disturbances follow a known probabilistic model, or robust control, which plans for the worst-case, online nonstochastic control operates in a more general and challenging setting.
In this framework, both the disturbances and cost functions are chosen by an adversary, meaning they may follow no statistical pattern and could even be designed to hinder the controller’s decisions at every step. As a result, the goal is no longer to compute a fixed optimal control policy but rather to learn and adapt dynamically in response to an evolving sequence of perturbations and costs.
A fundamental difficulty in control is the computational tractability of finding the best policy. Classical control methods often rely on dynamic programming, which, in general, suffers from the curse of dimensionality. Even in simple settings, computing an optimal policy can be intractable due to the complexity of solving constrained optimization problems over trajectories. As pointed out by Rockafellar, finding an optimal control policy in arbitrary convex cost settings can be computationally prohibitive, particularly when state and control constraints are involved.
To address this, online nonstochastic control shifts the focus from computing an optimal policy to competing with a benchmark class of policies using regret minimization. Instead of optimizing a predefined objective function, the controller seeks to minimize its total loss relative to the best policy that could have been chosen in hindsight. This approach bypasses the need to solve an intractable dynamic optimization problem upfront while still achieving performance guarantees over time.
The fundamental challenge in online nonstochastic control can be summarized as follows:
  • The controller does not have prior knowledge of the cost functions or disturbances.
  • At each time step, the controller must select a control input based only on past observations.
  • The performance of the controller is evaluated against a benchmark: the best fixed policy that could have been chosen in hindsight.
This formulation naturally connects to online learning and online convex optimization, where decisions must be made sequentially in uncertain environments. We now provide a formal definition of the problem.
@@
Consider a dynamical system evolving according to the equation: $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),$ where:
  • $\mathbf{x}_{t}\in\mathbb{R}^{d_{\mathbf{x}}}$ is the system state at time $t$,
  • $\mathbf{u}_{t}\in\mathbb{R}^{d_{\mathbf{u}}}$ is the control input at time $t$,
  • $\mathbf{w}_{t}\in\mathbb{R}^{d_{\mathbf{w}}}$ represents an external perturbation, which is chosen adversarially.
At each time step, the controller ${\mathcal{A}}$ selects $u_{t}$ based on past observations, then incurs a cost defined by an arbitrary convex function $c_{t}:\mathbb{R}^{d_{\mathbf{x}}}\times\mathbb{R}^{d_{\mathbf{u}}}\to\mathbb{R}$.
The objective is to design a control strategy that minimizes regret, defined as the difference between the total cost incurred by the controller and the best fixed policy in hindsight:
@@
where $\Pi$ is a benchmark class of policies, and $(x_{t}^{\pi},u_{t}^{\pi})$ represent the state and control sequence obtained under the best policy $\pi$ in hindsight.
Henceforth, if $T,\Pi$ and ${\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers. The goal of an online control algorithm is to achieve sublinear regret,
$\mathrm{{regret}}_{T}=o(T),$ (7.1)
which ensures that, as $T\to\infty$, the controller performs competitively with the best fixed policy.
Improper Learning and Convex Relaxation
A key technique that enables online nonstochastic control algorithms to remain computationally efficient is improper learning through convex relaxation. Instead of directly competing with a predefined class of policies, these algorithms leverage a different policy class that allows for efficient optimization while still achieving strong performance guarantees.
This approach is crucial because enforcing exact constraints on the policy space can lead to computational hardness results, as previously discussed. By relaxing the policy representation, for example using convex approximations or surrogate loss functions, controllers can sidestep these intractabilities while still competing effectively with the best policy in hindsight.
In the next section, we introduce algorithmic solutions that utilize this principle to achieve low-regret online control under adversarial conditions.